# Constrained Decoding for Robotics Foundation Models

### v3

@@ 4.1 Implementational Details
We comprehensively evaluate our constrained decoding framework on procedurally generated AI2-THOR [23] indoor scenes with diverse objects and layouts using three state-of-the-art (SOTA) RFMs: Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3] and Flare [4]. All three are large transformer-based embodied agents trained on extensive language-conditioned robot trajectory datasets. These models achieve strong zero-shot generalization for a vast variety of navigation tasks that span open vocabulary object-goal navigation (“find a mug”), room-to-room traversal (“visit all rooms”), waypoint-based navigation (“move three meters forward and stop near the red rug”), and attribute-conditioned variants (“locate the chair closest to the refrigerator in the kitchen”). These models also demonstrate reliable zero-shot transfer to real-world environments, achieving robust task satisfaction.
@@
In addition, these models capture three different training paradigms for RFMs. SPOC is trained purely with imitation learning from shortest-path rollouts. Poliformer employs a hybrid approach that combines reinforcement learning and imitation learning, enabling it to learn long-horizon structure while retaining expert priors. Flare adopts a large-scale pretraining plus fine-tuning on embodied navigation data in line with recent foundation model training paradigms. This diversity in training paradigms allows us to evaluate the applicability of SafeDec across different learning regimes.
@@
In this work, we address safety specifications for robotics and, therefore, select those most relevant to real-world deployment. In particular, we focus on an important class of safety specifications called invariants, which are specifications that must be enforced at every reachable state of the system (e.g., “always avoid an unsafe region”). We enforce geofencing and obstacle avoidance by encoding them as invariant specifications in STL. Specifically, we generate random regions in the configuration space that the robot must either avoid (obstacle zones) or remain within (safe zones), and apply these constraints in real time during execution. The specifications used are: $\varphi_{\text{geofence}}=\mathbf{G}\!\left(\bigvee_{i=1}^{N}\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$, $\varphi_{\text{avoid}}\;=\;\mathbf{G}\!\left(\bigwedge_{i=1}^{N}\neg\Bigl(x_{i}^{\mathrm{L}}\leq x\leq x_{i}^{\mathrm{U}}\;\wedge\;z_{i}^{\mathrm{L}}\leq z\leq z_{i}^{\mathrm{U}}\Bigr)\right)$. The size of the regions for $\varphi_{avoid}$ is 1 $m^{2}$. For $\varphi_{geofence}$, we randomly pick a subset of rooms in each house and use each chosen room’s full bounds. These safety specifications are common for robot learning applications [24, 25, 26, 27]. We encode our test STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel [14]. This ensures minimal inference overhead at runtime ($10^{-5}$ s per timestep) , which is crucial for policy deployment. For our dynamics model, we assume a unicycle model, an approximate first-order dynamics abstraction widely used in the robotics literature for analysis and control [28]. This representation captures the essential kinematics of motion in the plane and is widely used because it is applicable for diverse robotic platforms.
@@ 4.2 Experimental setup
We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism. The filtering mechanism picks a default action (turning left or right in place) upon predicted violation of the safety specification, similar to the Simplex architecture [29]. Simplex architecture is a classic scheme in which a high-performance advanced controller is continuously monitored by a provably safe but less capable backup controller. Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons. We evaluate performance using two main metrics: STL Satisfaction Rate (STL St), defined as the proportion of trajectories that satisfy the specified STL formula, and Task Success Rate (SR), which measures standard task success. The three main research questions we investigated in this paper:
  1. 1.
    RQ1: Do HCD and RCD provide higher STL satisfaction than the unconstrained baselines?
  2. 2.
    RQ2: Do HCD and RCD preserve task success rates comparable to the unconstrained baselines?
  3. 3.
    RQ3: Does RCD achieve better task success than HCD while maintaining high STL satisfaction?
$\bm{\phi}_{avoid}$: STL / SR (% $\uparrow$) $\bm{\phi}_{geofence}$: STL / SR (% $\uparrow$)
Decoding SPOC Flare PoliFormer SPOC Flare PoliFormer
Unconstrained 72.0 / 82.5 75.5 / 82.0 77.0 / 82.5 78.0 / 81.5 68.0 / 81.0 73.0 / 81.5
Filtering 100.0 / 72.0 100.0 / 78.5 100.0 / 75.5 100.0 / 72.0 100.0 / 66.5 100.0 / 67.5
HCD 100.0 / 72.5 100.0 / 81.0 100.0 / 78.5 100.0 / 76.5 100.0 / 67.5 100.0 / 72.5
RCD 93.0 / 76.0 83.0 / 82.5 87.5 / 83.5 95.5 / 80.0 80.0 / 71.5 85.5 / 77.5
Table 1: Comparison by decoding technique across models (SPOC, FLARE, PoliFormer) for specifications $\phi_{avoid}$ and $\phi_{geofence}$. Each cell reports STL satisfaction / success rate (%). Higher is better (↑).
(a) Unconstrained v/s HCD
(b) Unconstrained v/s RCD
Figure 2: Qualitative comparison of decoded trajectories for a sample scene. Each plot shows a top down view of an overlay of trajectories starting from the white dot under the instruction “find an alarm clock”. The unconstrained model passes through two forbidden regions (red squares) on the way to the target object located on the table. In contrast, HCD (left) and RCD (right) modify the trajectories to respect STL safety specifications while still reaching the goal. More visualizations are available in the Appendix A.2.
@@ 4.3 Results
Our results are highlighted in Table 1. We also visualize sample trajectories in Figure 2 for one scene and task. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
RQ1 – STL satisfaction.
[…]
RQ2 – Task completion.
@@
Simplex-style filtering attains high satisfaction but sacrifices task success because the agent takes predefined safe actions. HCD shows similar behavior: although safety is maximized, success rates are consistently 5–10% lower than the baseline across models and specifications. However, as HCD factors in base model logits, it is able to achieve higher task satisfaction compared to Simplex-style filtering. In contrast, RCD preserves success rates much closer to the unconstrained level. For $\phi_{avoid}$, RCD achieves 82–85% success compared to 82–83% for the unconstrained controllers for Flare and PoliFormer. For $\phi_{geofence}$, it maintains 77–80% compared to 81–82% unconstrained for SPOC and PoliFormer. However, we note that RCD does not fully recover success in every case: on SPOC with $\phi_{avoid}$ and Flare with $\phi_{geofence}$, task success remains several points below the unconstrained baseline. Nevertheless, RCD enforces safety while avoiding the large performance penalty observed with filtering.
RQ3 – RCD vs. HCD.
@@
While both HCD and RCD improve safety over the unconstrained baseline, they differ in how they balance constraint satisfaction with task success. HCD enforces strict STL satisfaction that results in frequent conservatism and lower successful task completion rates. In contrast, RCD’s soft penalization leads to higher task success while still maintaining reasonable STL satisfaction. These results show that RCD achieves a better trade-off between safety and goal-directed behavior, especially in settings where occasional low-risk actions can lead to higher long-term rewards.
@@
Overall, Our proposed techniques effectively enforce safety STL specifications during policy execution. HCD ensures full compliance, but occasionally sacrifices task success due to strict truncation. RCD strikes a balance, offering high satisfaction rates and robust performance. This highlights the feasibility of combining learning-based models with formal safety constraints.
@@ 4.4 Ablation studies
Does inaccurate dynamics modeling substantially reduce STL satisfaction? In this work, we assume a simple unicycle dynamics model due it’s generalization capability for diverse robotic platforms. Although this represents a high-level abstraction of true dynamics, such modeling simplifications are standard in the formally assured robot safety literature [28]. However, both RCD and HCD depend on this assumption and inaccurate modeling can impact STL satisfaction. To evaluate the impact of inaccurate dynamics modeling, we conducted an ablation in which we inject gaussian perturbations into the dynamics (0.01 m per step translational noise i.e. 5% of nominal forward step, 1 ° per step rotational noise i.e. 3.3% of yaw step) for both HCD and RCD. Our results are visualized in Figure 3. Across all three base models, the drop in STL satisfaction rates from baseline to noisy dynamics is relatively small. We observe that SafeDec shows graceful degradation under significant per-step disturbances.
@@
How do varying values of $\beta$ affect success rate and STL satisfaction in RCD? As highlighted in section 3.3, RCD uses a hyperparameter $\beta$ that modulates the trade-off between specification adherence and the original task objective. The value of $\beta$ affects the relative weighting between robustness and base logits. To investigate the impact of $\beta$ on the success rate and STL satisfaction, we performed an ablation with varying values for $\beta$ for Flare and PoliFormer. We observe that as $\beta$ increases for PoliFormer, both STL satisfaction and success rate improve in tandem until $\beta$ = 10, suggesting that moderate regularization can actually aid policy execution. Beyond this, STL satisfaction continues to improve but at the cost of lower success rates. For Flare, larger $\beta$ values improve STL satisfaction but reduce success rates. These results highlight that the influence of $\beta$ is model-dependent but in general demonstrate that SafeDec provides a tunable mechanism to balance safety and performance objectives.
@@ 5 Related Work
Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [30], safe reinforcement learning [31], temporal logic-based shielding and synthesis approaches [32, 33, 34]. Recently, with the advent of vision language action models and their impressive generalizable capabilities for manipulation, navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models. Although classical methods offer formal guarantees, they either require pretraining/fine-tuning stage interventions or designing a new classical controller for each safety specification, which can be restrictive. For example, SafeVLA [35] fine-tunes pre-trained foundation models with task-specific safety costs, achieving strong performance in Safety-CHORES tasks. However, the safety specification is expected to be embedded in the training data and loss, meaning the model cannot generalize to new safety constraints at test time. In contrast, ASIMOV [36] explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees. Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model. The closest to our work is SELP [37] that proposes LTL-constrained decoding for language model-based plan generation. However, SELP is unsuitable for STL because its Boolean predicate-based LTL cannot encode numeric bounds (e.g. $x–x_{goal}||<0.1m$) and does not possess quantitative semantics, which is crucial for ranking actions. These techniques are also tailored to high-level plans from LLMs, not to the per-step low-level actions generated online by policies.
@@ 6 Limitations and Future Work
In this work, we introduce a constrained decoding framework for enforcing safety specifications for large transformer based robot policies. Our approach enables runtime adaptation to novel safety specifications without retraining. Through experiments across multiple simulated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates. Our approach makes two critical assumptions that can be a limiting factor. First, we assume access to specifications that are defined over the state space and that these specifications are generated by roboticists. Although this is a common situation for safety critical deployment contexts like aerial robotics [38, 39], these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation. We plan to overcome this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic [40] and using large language models for generating high level specifications automatically [41]. Second, our approach also assumes access to an approximate dynamics model to evaluate the impact of actions on future trajectories. While a common assumption for provably safe robotics [28], This can limit applicability of our approach. However, it is possible to mitigate this via learned dynamics models such as MoSim [42] or world models proposed in Zhou et al. [43], Micheli et al. [44] and we plan to explore this in future work.

### v4

@@ 4.1 Implementational Details
We comprehensively evaluate our constrained decoding framework on procedurally generated AI2-THOR (Kolve et al., 2022) indoor scenes with diverse objects and layouts. AI2-THOR is a photorealistic 3D simulation environment built on the Unity engine and is a widely adopted benchmark and training environment for embodied navigation research. We use three state-of-the-art (SOTA) generalist robot navigation policies: Shortest Path Oracle Clone (SPOC) (Ehsani et al., 2024), PoliFormer (Zeng et al., 2025) and Flare (Hu et al., 2025). All three are large transformer-based policies trained on extensive language-conditioned robot trajectory datasets. These models achieve strong zero-shot generalization for a vast variety of navigation tasks that span open vocabulary object-goal navigation (“find a mug”), room-to-room traversal (“visit all rooms”), waypoint-based navigation (“move three meters forward and stop near the red rug”), and attribute-conditioned variants (“locate the chair closest to the refrigerator in the kitchen”). These models also demonstrate reliable zero-shot transfer to real-world environments, achieving robust task satisfaction.
@@
In addition, these models capture three different training paradigms for generalist robot navigation policies. SPOC is trained purely with imitation learning from shortest-path rollouts. Poliformer employs a hybrid approach that combines reinforcement learning and imitation learning, enabling it to learn long-horizon structure while retaining expert priors. Flare adopts a large-scale pretraining plus fine-tuning on embodied navigation data in line with recent foundation model training paradigms. This diversity in training paradigms allows us to evaluate the applicability of SafeDec across different learning regimes.
In this work, we address safety specifications for robotics and, therefore, select those most relevant to real-world deployment. Our specifications are highlighted in Table 1. These specifications are inspired by common safe navigation specifications as highlighted in Menghi et al. (2019) and have varying levels of complexity. These specifications capture state-based invariants ($\phi_{avoid},\phi_{geofence}$), history-dependent conditionals ($\phi_{no\_return}$) and temporal ordering constraints ($\phi_{ordered}$).
Table 2: Comparison by decoding technique across models (SPOC, Flare, PoliFormer) for specifications $\phi_{avoid}$, $\phi_{ordered}$, and $\phi_{no\_return}$. Each cell reports STL satisfaction / success rate (%). Higher is better (↑).
$\bm{\phi}_{avoid}$: STL St / SR (% $\uparrow$) $\bm{\phi}_{ordered}$: STL St / SR (% $\uparrow$) $\bm{\phi}_{no\_return}$: STL St / SR (% $\uparrow$)
Decoding SPOC Flare PoliFormer SPOC Flare PoliFormer SPOC Flare PoliFormer
Unconstrained 72.0 / 82.5 75.5 / 82.0 77.0 / 82.5 76.0 / 80.0 72.0 / 80.5 71.0 / 81.5 84.0 / 79.5 82.0 / 81.0 82.0 / 79.0
Filtering 100.0 / 72.0 100.0 / 78.5 100.0 / 75.5 100.0 / 69.0 100.0 / 70.0 100.0 / 67.5 100.0 / 68.5 100.0 / 74.0 100.0 / 66.5
HCD 100.0 / 72.5 100.0 / 81.0 100.0 / 78.5 100.0 / 70.5 100.0 / 73.0 100.0 / 72.5 100.0 / 71.0 100.0 / 76.5 100.0 / 69.5
RCD 93.0 / 76.0 83.0 / 82.5 87.5 / 83.5 91.0 / 75.0 89.0 / 79.5 95.0 / 78.0 96.0 / 77.5 96.0 / 80.0 94.0 / 74.0
@@
We encode our test STL specifications using an efficient computational graph-based STL library called STLCG++ that can evaluate multiple state signals in parallel (Kapoor et al., 2025c). This ensures minimal inference overhead at runtime ($10^{-5}$ s per timestep) , which is crucial for policy deployment. For our specifications, we generate random regions in the configuration space at runtime and use their coordinates as spatial predicates. For our dynamics model, we assume a unicycle model, an approximate first-order dynamics abstraction widely used in the robotics literature for analysis and control (Cohen et al., 2024). This representation captures the essential kinematics of motion in the plane and is widely used because it is applicable for diverse robotic platforms.
@@ 4.2 Experimental setup
We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism. The filtering mechanism picks default actions that cannot immediately violate the specification (eg. turning in place) upon predicted violation of the safety specification, similar to the Simplex architecture (Sha, 2001). Simplex architecture is a classic scheme in which a high-performance advanced controller is continuously monitored by a provably safe but less capable backup controller. Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons. Additionally, our base policies represent actions as a discrete set of low-level action primitives (e.g., move forward/backward, lateral strafing, rotation, end, and a few other discrete navigation actions exposed by each policy). At each timestep, the policy outputs a categorical distribution over this shared action set. We evaluate performance using two main metrics: STL Satisfaction Rate (STL St), defined as the proportion of trajectories that satisfy the specified STL formula, and Task Success Rate (SR), which measures standard task success.
@@ 4.3 Results
Our results are highlighted in Table 2. For brevity, we provide the results for $\phi_{geofence}$ in 4 since it is syntactically similar to $\phi_{avoid}$. We also visualize sample trajectories in Figure 2 for one scene and task. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
Do HCD and RCD provide higher STL satisfaction than the unconstrained baselines? Both HCD and RCD consistently improve STL satisfaction relative to the unconstrained baselines across all models and specifications. All policies with RCD show significant gains, with improvements ranging from 8-24 percentage points. HCD achieves perfect STL satisfaction, with improvements ranging from 16-29 percentage points. We observe that the Simplex-style filtering baseline achieves similar STL satisfaction rate as HCD. This parity is expected, as both methods block any action predicted to violate the specification.
Figure 2: Qualitative comparison (Unconstrained vs. HCD). Top-down visualization of trajectories for a sample scene under the instruction “find the sofa”. The unconstrained policy passes through forbidden regions (red squares), while Hard-Constrained Decoding (HCD) enforces STL safety specifications and generates trajectories that avoid unsafe regions while still reaching the sofa (center-left). Additional visualizations are provided in Appendix A.2.
Which safety enforcement technique achieves task success rates comparable to the unconstrained baselines?
@@
As highlighted in Table 2, Simplex-style filtering attains high STL satisfaction but sacrifices task success because the agent takes predefined safe actions. HCD shows similar behavior: although safety is maximized, success rates are consistently lower than the unconstrained baseline across models and specifications. However, as HCD factors in base model logits, it is able to achieve higher task satisfaction compared to Simplex-style filtering with improvements ranging from 0.5-3 percentage points. In contrast, RCD preserves success rates much closer to the unconstrained level with success-rate drops of 0.5–6.5 percentage points (smaller drop is better). However, we note that RCD does not fully recover success for certain models and specifications, task success remains several points below the unconstrained baseline. Nevertheless, RCD enforces safety while avoiding the large performance penalty observed with filtering.
@@
Does RCD achieve better task success than HCD while maintaining high STL satisfaction? While both HCD and RCD improve safety over the unconstrained baseline, they differ in how they balance constraint satisfaction with task success. HCD enforces strict STL satisfaction that results in frequent conservatism and lower successful task completion rates. In contrast, RCD’s soft penalization leads to higher task success while still maintaining reasonable STL satisfaction. These results show that RCD achieves a better trade-off between safety and goal-directed behavior, especially in settings where occasional low-risk actions can lead to higher long-term rewards.
@@
Overall, Our proposed techniques effectively enforce safety STL specifications during policy execution. HCD ensures full compliance, but occasionally sacrifices task success due to strict truncation. RCD strikes a balance, offering high satisfaction rates and robust performance.
4.4 Comparison with safe reinforcement learning baseline
Table 3: Comparison with SafeVLA on the SafetyChores benchmark. Lower cost indicates fewer safety violations, while higher success denotes better task completion.
Method Cost $\downarrow$ SR (%) $\uparrow$
ISA (SafeVLA) 0.205 86.5
ISA + HCD 0.015 82.0
ISA + RCD 0.060 86.0
FLARE 0.192 82.0
FLARE + HCD 0.115 84.0
FLARE + RCD 0.155 79.0
We additionally compare SafeDec with SafeVLA (Zhang et al., 2025), a safe reinforcement learning method that proposes an integrated safety approach (ISA) using explicit safety costs. To enable a fair comparison, we focus on the fragile-object and dangerous-object specifications in the SafetyChores benchmark (Zhang et al., 2025), which closely align with our setting, and construct corresponding STL specifications for evaluation. We report the composite safety cost defined in SafeVLA, which directly reflects the degree of safety violations across both object classes, along with task success.
Our results are highlighted in Table 3. Our findings indicate that both HCD and RCD substantially reduce safety violations relative to their respective vanilla baselines while largely preserving task success. In particular, integrating HCD with ISA (SafeVLA) achieves an order-of-magnitude reduction in safety cost (0.205 → 0.015), demonstrating effective enforcement of fragile and dangerous object specifications with minimal impact on task performance. We also find that RCD+ISA preserves task success close to vanilla ISA while reducing safety violations. Our results show that SafeDec, despite being a training-free method, preserves task satisfaction close to the safeRL baseline (ISA) for our specifications. Additionally, combining SafeDec with the ISA policy leads to further reduction in safety costs, highlighting its complementary benefits to data-driven safe learning.
@@ 4.5 Ablation studies
Does inaccurate dynamics modeling substantially reduce STL satisfaction? In this work, we assume a simple unicycle dynamics model due it’s generalization capability for diverse robotic platforms. Although this represents a high-level abstraction of true dynamics, such modeling simplifications are standard in the formally assured robot safety literature (Cohen et al., 2024). However, both RCD and HCD depend on this assumption and inaccurate modeling can impact STL satisfaction. To evaluate the impact of inaccurate dynamics modeling, we conducted an ablation in which we inject gaussian perturbations into the dynamics (0.01 m per step translational noise i.e. 5% of nominal forward step, 1 ° per step rotational noise i.e. 3.3% of yaw step) for both HCD and RCD. Our results are visualized in Figure 3(a). Across all three base models, the drop in STL satisfaction rates from baseline to noisy dynamics is relatively small. We observe that SafeDec shows graceful degradation under significant per-step disturbances.
@@
How do varying values of $\beta$ affect success rate and STL satisfaction in RCD? As highlighted in section 3.3, RCD uses a hyperparameter $\beta$ that modulates the trade-off between specification adherence and the original task objective. The value of $\beta$ affects the relative weighting between robustness and base logits. To investigate the impact of $\beta$ on the success rate and STL satisfaction, we performed an ablation with varying values for $\beta$ for Flare and PoliFormer. As visualized in Figure 3(b), We observe that as $\beta$ increases for PoliFormer, both STL satisfaction and success rate improve in tandem until $\beta$ = 10, suggesting that moderate regularization can actually aid policy execution. Beyond this, STL satisfaction continues to improve but at the cost of lower success rates. For Flare, larger $\beta$ values improve STL satisfaction but reduce success rates. These results highlight that the influence of $\beta$ is model-dependent but in general demonstrate that SafeDec provides a tunable mechanism to balance safety and performance objectives.
@@ 5 Related Work
Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) (Ames et al., 2019), safe reinforcement learning (Gu et al., 2024), and temporal logic-based shielding approaches (Alshiekh et al., 2017). Recently, with the advent of vision language action models and their impressive generalizable capabilities for embodied navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models. Although classical methods offer formal guarantees, they either require training stage interventions or designing a new classical controller for each safety specification, which can be restrictive. For example, SafeVLA (Zhang et al., 2025) uses safeRL with task-specific safety costs, achieving strong performance in SafetyCHORES tasks. However, the safety specification is expected to be embedded in the training data and loss, meaning that the model cannot generalize to new safety constraints at test time. In contrast, ASIMOV (Sermanet et al., 2025) explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees. Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model. There has been some work on logic constrained decoding such as SELP (Wu et al., 2025) that proposes LTL-constrained decoding for language model-based plan generation. However, SELP is unsuitable for STL because its Boolean predicate-based LTL cannot encode numeric bounds (e.g. $x–x_{goal}||<0.1m$) and does not possess quantitative semantics, which is crucial for ranking actions. Moreover, SELP operates at the symbolic planning level, constraining high-level plans produced by LLMs offline, whereas our method enforces safety over per-step low-level actions generated online by embodied policies.
@@ 6 Limitations and Future Work
In this work, we introduce a constrained decoding framework for enforcing safety specifications for large transformer based robot navigation policies. Our approach enables runtime adaptation to novel safety specifications without retraining. Through experiments across multiple procedurally generated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates. Our approach makes two critical assumptions that can be a limiting factor. First, we assume access to specifications that are defined over the state space and that these specifications are generated by roboticists. Although this is a common situation for safety critical deployment contexts like aerial robotics (Aloor et al., 2023; Luckcuck et al., 2019), these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation. We plan to overcome this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic (Kapoor et al., 2025a) and using large language models for generating high level specifications automatically (Li et al., 2025). Second, our approach also assumes access to an approximate dynamics model to evaluate the impact of actions on future trajectories. While a common assumption for provably safe robotics (Cohen et al., 2024), this can limit applicability of our approach. However, it is possible to mitigate this via learned dynamics models such as MoSim (Hao et al., 2025) or world models proposed in Zhou et al. (2025); Micheli et al. (2023) and we plan to explore this in future work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.